The yield on 1-year Treasury securities is 6%, 2-year securities yield 6.2%, 3-year securities yield 6.3%, and 4-year securities yield 6.5%. There is no maturity risk premium. Using expectations theory and geometric averages, forecast the yields on the following securities:

A 1-year security, 1 year from now?

A 1-year security, 2 years from now?

A 2-year security, 1 year from now?

A 3-year security, 1 year from now?
Answer #1

As per expectation theory, long term interest rate is simple a weighted average of current and expected future short-term interest rate provided MRP is zero.

1)

(1 + two-year yield)2 = (1 + one-year yield) x (1+ future one-year yield)

(1+ 0.062) 2 = (1 + 0.06) x (1+ future one-year yield)

(1+ 0.062) 2 / (1 + 0.06) = 1+ future one-year yield

1+ future one-year yield = (1.062) 2/ (1.06)

                                           = 1.127844/1.06

                                          = 1.06400377358491

Future one-year yield = 1.06400377358491 – 1 = 0.06400377358491 or 6.4 %

2)

(1 + three-year yield)3 = (1 + two-year yield)2 x (1+ future one-year yield)

(1 + 0.063)3 = (1 + 0.062) x (1+ future one-year yield)

(1.063)3 = (1.062) 2 x (1+ future one-year yield)

(1.063)3 / (1.062) 2 = (1+ future one-year yield)

1+ future one-year yield = 1.201157047/1.127844

                                               = 1.0650028257454

Future one-year yield = 1.0650028257454 – 1 = 0.0650028257454 or 6.5 %

3)

(1 + three-year yield)3 = (1 + one-year yield) x (1+ future two-year yield)2

(1 +0.063)3 = (1 + 0.06) x (1+ future two-year yield)2

(1.063)3 = (1.06) x (1+ future two-year yield)2

(1.063)3/ (1.06) = (1+ future two-year yield)2

(1+ future two-year yield)2 = 1.201157047/1.06

                                               = 1.1331670254717

Future two-year yield = √ (1.1331670254717) – 1

                                        = 1.06450318246199 – 1

                                        = 0.06450318246199 or 6.45 %

4)

(1 + four-year yield)4 = (1+ one-year yield) x (1 + future three-year yield)3

(1 +0.065)4 = (1 + 0.06) x (1 + future three-year yield)3

(1.065)4 = (1.06) x (1 + future three-year yield)3

(1.065)4 / (1.06) = (1 + future three-year yield)3

(1 + future three-year yield)3 = 1.286466350625/1.06

                                               = 1.21364750058962

1 + future three-year yield = 3 √ 1.21364750058962 = 1.06667190227378

Future three-year yield =1.06667190227378 -1 = 0.06667190227378 or 6.67 %